OpenStudy (anonymous):

Which of the following is a polynomial with roots - square root of 3, square root of 3, and -2 ? Choose one answer. a. x3 - 2x2 - 3x + 6 b. x3 + 2x2 - 3x - 6 c. x3 - 3x2 - 5x + 15 d. x3 + 3x2 - 5x - 15

OpenStudy (auctoratrox):

b. just take (x - sqrt3)(x + sqrt3)(x + 2) = ...

jhonyy9 (jhonyy9):

(x-sqrt3)(x+sqr3)(x+2)=(x^2 -3)(x+2)=x^3 +2x^2 -3x-6 - so than will be b.

OpenStudy (anonymous):

Thanks so much!